OpenStudy (anonymous):

What is the value of x? Round your answer to the nearest tenth if necessary.

OpenStudy (studygurl14):

what's the equation and what did you get?

OpenStudy (anonymous):



OpenStudy (alexandervonhumboldt2):

x^2=7^2+4^2 by pythgorian thorem

OpenStudy (alexandervonhumboldt2):

x^2=65 x=sqrt565=use calc and round

OpenStudy (alexandervonhumboldt2):

sqrt65 sorry

OpenStudy (anonymous):

I GOT 65 am i correct sorry 4 caps

OpenStudy (alexandervonhumboldt2):

no not correct

OpenStudy (alexandervonhumboldt2):

to need to sqrt65

OpenStudy (alexandervonhumboldt2):

8.0622577482985496523666132303038=round is to tenth place

OpenStudy (anonymous):

8.06

OpenStudy (alexandervonhumboldt2):

nono tenth place so 8.1

OpenStudy (anonymous):

o sorry about that so its 8.1

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

its 8.1?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

K thanks